# Development first Performance evaluation of multi element monolithic HPGe detector for spectroscopy applications

### v1

@@ Abstract
We present the first operational prototype of a high-purity Germanium (HPGe) detector developed within the framework of the European LEAPS-INNOV project. This prototype features a monolithic, multi-element sensor optimized for high-resolution X-ray spectroscopy in the hard X-ray regime. The study includes a comprehensive laboratory-based characterization of the detector’s spectral and spatial performance, and an on-beam evaluation at the BM05 beamline of the ESRF synchrotron facility using monochromatic X-rays in the 20–50 keV energy range. Detailed characterization results are reported, including sensor performance metrics such as energy resolution, count rate, and spatial response uniformity.
@@ 1 Physics Motivation
Fundamentally, the choice of Germanium over Silicon is motivated by its higher atomic number (Z=32) vs. (Z=14) for Si, which enhances the photoelectric absorption cross-section, particularly in the hard X-ray regime, meaning that HPGe detectors exhibit substantially higher quantum efficiency for photon energies above 20 keV [MOHANTY2008186], allowing for more effective detection of weak signals in dilute samples containing heavy elements [Cole2001RareEarthEXAFS]. Additionally, Germanium’s smaller bandgap (0.67 eV) enables finer energy resolution by generating more charge carriers per absorbed photon, improving the resultant signal to background/noise ratio (SNR) [FARROW1995567, dennis2019].
@@
The monolithic design avoids boundaries between separate crystals, which can cause signal variations due to gaps or misalignment. Using high-purity germanium reduces the number of crystal defects, thereby reducing charge-carrier trapping. As a result, the detector can achieve high energy resolution and consistent signal response across the entire active area [Hexgonal_pad_N_tartoni], attaining qualities critical in spectroscopy applications.
@@
Great care is demanded while fabricating such detectors, as they require ultra-high purity and operate in a cryogenic environment (close to 77 ) to suppress thermal noise.
@@
Building on this motivation, Work Package 2 of the European LEAPS-INNOV [LEAPSINNOV2020] project has taken a step towards developing a new generation of HPGe detectors [F_ORSINI_FIRSTARTICLE, Goyal_2025, Gimenez_2025]. We have developed two such detectors with 10-element monolithic sensors, featuring pixel areas of 5 mm2 and 20 mm2. These detectors offer the capability of handling high input count rates (20-250 kcps/mm2) in a broad energy range of (20- 100 keV), while maintaining a reliable energy resolution aiming 180 eV FWHM at 5.9 keV, within the integration time window of 1 µsec [Goyal:2025oph]. In addition, the Digital Pulse Processing technique (DPP) will further improve detector performance by electronically rejecting charge-sharing events, degrading the resultant spectral resolution [QuantumXspress3] due to incomplete charge collection.
@@
This article presents the full development and characterization of the first HPGe prototype. In section 2, we describe the detector design, including the monolithic design of the sensor, the cryogenic cooling, mechanical architecture, and TETRA ASIC-based new front-end electronics. Section 3 discusses the laboratory-based characterization with detailed analysis of signal rise time, Power spectral density, fluorescence measurements with metallic foil samples, and surface scanning using a micro spot beam. In section 4, we discuss the first on-beam evaluation of the detector, followed by a detailed analysis of defects observed in several pixels. We provide a phenomenological model to derive the depth of these defects. And finally, we conclude with a summary of detector performance so far, and future perspectives for improving the current performance in future prototype versions.
@@ 2 The XAFS-DET germanium detector
The XAFS-DET detector, shown in Fig. 1, is composed of a compact detector head, a detector body, and ancillary elements. Other synchrotron spectroscopy detectors typically follow the same schematic structure. The detector head, which includes the germanium sensor, the front-end electronics, and their interconnections, is under vacuum at cold temperature (77 ). The detector body comprises the cooling system, the cryostat with its flanges, and the back-end electronics. Ancillary elements include the digital pulse processor, voltage power supplies, and equipment control and monitoring.
@@ 2.1 Sensor Design
The sensor is a multi-element high-purity germanium sensor manufactured by Mirion Technologies (CANBERRA). A scheme of the sensor is shown in Fig. 2. It has a square shape (20 mm length) and a thickness of 4.1 mm. The sensor is a p-type Ge bulk working in hole-collection mode. The front side (facing the incident X-rays) is phosphor-implanted (n-type doped). In contrast, the back side is boron-implanted (p-type doped) and is segmented into a central hexagonal pixel, six trapezoidal pixels situated in a ring around the central one, and three outer ring-shaped pixels, which serve as ancillary pixels to reject charge-sharing events by sharing the incident charge with the guard ring. The dimensions of hexagonal and trapezoidal pixels have been defined to achieve an area of 20 mm2 per pixel. Aluminium contacts were evaporated on both sides, and the inter-pixel spacing provides an insulation of 10 T. A positive potential difference is established by applying a high voltage to the frontside window contact via a high-voltage Rogers PCB, while backside pixels are connected to front-end board pads via copper-gold-plated pogo pin connectors (Model P706 from Peak Test, 6.9 mm long with 0.2 mm-diameter tips). The full depletion voltage is around +33 V, and the detector was operated at +75 to +300 V during characterization. We measured the leakage current of each pixel to be less than 1 pA at liquid nitrogen temperature.
@@ 2.2 Electronics
TETRA ASIC is a four-channel charge-sensitive (CSA) preamplifier working in hole collection mode with synchronous reset [6154396]. The preamplifier gain setting (low, medium, or high), reset features (threshold, width, and external trigger), and channel disabling can be configured by the user via a 32-bit shift register. All four channels share a test-pulse input to characterize signal fall time and crosstalk between channels without the sensor, as well as to simulate input leakage current [CMOS_CSA]. The FEB is a circular board (20 mm radius, 1.4 mm thick) made of ceramics, as shown in Fig. 3 (left). It links the ten pogo pin pads (0.4 mm radius), on its front side, to ten input channels of three TETRA ASICs (two with three channels enabled, one with all four channels enabled) located on its back side. The FEB back side also houses a thermistor. Output signals and TETRA power supplies are routed to the BEB via a 60 cm-long flex kapton PCB cable with a subD25 connector. The cable is glued to the FEB back side surface, and electrical continuity is made by wire bonding, while the subD25 connector is screwed to the back flange feedthrough of the vacuum cube.
@@
The BEB is a set of three PCB boards (buffer, bias, and control) designed to bias and configure three TETRA ASICs and buffer 10 analog channels. A conceptual schema is shown in Fig. 3 (right). The buffer board is directly connected to the subd25 feedthrough to minimize the path of ASIC outputs. The ten signals are buffered and routed to ten SMA connectors, which are connected to the Digital Pulse Processor via coaxial RF cables. The bias board provides power to itself and the other two boards, applies the voltage-level shifting of the ASIC program bus, and acquires the FEB thermistor signal. The bias board also includes three SMA connectors: two inputs for the external reset pulse and the test pulse, and one output for the inhibit reset signal. The control board monitors and controls BEB features via an FPGA unit and communicates with the user via USB 3.0 or Ethernet.
@@
The electronics performance is summarized in Table 1. The electronics step time evolution function without the sensor was characterized at room temperature(23-25°C) in terms of signal gain and fall-time (10% to 90% of the amplitude) for the three gain settings and for all channels. These features were measured from the mean waveform generated by a squared-test pulse (1 Volt amplitude, 10 ns rising edge), as illustrated in Fig. 4(a). Fall-time values of 19-23 ns were measured for all gains, and we verified that the gain ratios high-medium and medium-low were 2.0 and 3.0, see Fig. 4(b). We also quantified the absolute gain values later during lab tests with the sensor. In a second step, the range of values for reset duration (0.7-5.9 s) and threshold (-2.3-1.3 V) was measured. The reset period was longer than 4-5 sec, for a negligible leakage current. The period was shorter (around 23, 41, and 116 ms for high, medium, and low gain settings, respectively) with the leakage emulator enabled. Finally, the electronics noise (ENC) for peaking times between 0.1 and 10 s was measured, yielding values of 36-45 and 29-35 electrons for 1.0 and 10 s, respectively, within specifications. This series of tests was repeated at cold temperature (113 K), as discussed in [Goyal:2025oph].
@@
| ENC @ 1 µs Peaking Time | < 45 electrons | 36-45 electrons |
| ENC @ 10 µs Peaking Time | < 35 electrons | 29-35 electrons |
@@ 2.3 Cooling and mechanical design
The aluminum container, a copper arm holder, a horizontal copper cooling pipe, a thermal strap, and a vertical copper braid to the cryocooler evacuate the dissipated power from the front-end board [Quispe:2024wis].The pads of Indium foils link each interface. The cooling pipe is supported by two PEEK arm holders inside the inner cylinder, with small supports to minimize thermal losses. Two thermometers, glued to the vertical copper braid and the copper arm, monitor the temperature. The second one works as a feedback device for the cryocooler system. A zeolite box is attached to the vertical copper braid piece, which maintains vacuum integrity in static vacuum by adsorbing residual gases using pre-activated zeolite treated under vacuum at high temperatures. The box is equipped with a heating element and a thermometer, which are used to regenerate zeolites once they are saturated. The cooling chain of the detector is showcased using Fig.5 .
The cryocooler is mounted on a vacuum cube that holds the detector head in a front-facing tube. The cube’s three CF100 flanges provide dedicated feedthroughs for electronics (SubD25), cooling thermometers (SubD9), and HV bias. A separate flange houses the pumping port, vacuum gauge, and a SubD9 for the zeolite heater. A low-pass RC filter with a 6.5 Hz cut-off frequency conditions the HV signal to the sensor. A turbo pumping system achieves mbar vacuum, with residual gas analysis (MKS model) confirming only trace amounts of air and water.
Induced vibrations at the sensor head were measured using an Active Vibration Cancellation (AVC) system (model VSM101) produced by ifm Electronic. Acceleration values of 0.04-0.07 m/s2 at 1-2 frequencies above 50 Hz were measured in the three axes, which are within specifications.
@@ 3 Detector characterization in the laboratory
The procedures and results for each test are detailed in the following subsection.
@@ 3.1 Mean Signal risetime and preamplifier gain
The signal rise time and preamplifier gain were characterized using a 55Fe radioactive source emitting 5.9 and 6.4 keV X-rays. We operate the detector under low-flux conditions (approximately 10 kcps), and 100 K individual signal waveforms were acquired using a Lecroy digital oscilloscope operating at 0.5 Gs/s. We performed the measurements with bias voltages varying from +50 V to +330 V in 10 V increments.
Each recorded waveform was processed to extract the signal amplitude and rise time. The distribution of these parameters is analyzed as a function of bias voltage. The mean pulse amplitude was used to determine the preamplifier gain, yielding values between 1.99 and 2.06 mV/keV (0.02 mV/keV systematic error), which are within design specifications.
As shown in Fig. 6, the rise time distributions shift progressively toward shorter values as the bias voltage increases. This trend reflects the expected enhancement in hole drift velocity with stronger electric fields, reaching a minimum rise time of about 42 ns in the 280–330 V range.
@@ 3.2 Power Spectral Density Measurements
We measured the PSD over the frequency range of 100 Hz to 50 MHz using a fast digitizer (Model DT9862, by Data Translation), triggered 100 ms after the RESET signal to capture noise during the signal-processing window. The measurements were performed on the central pixel for all available TETRA ASIC gain settings. As shown in Fig. 7(a), the noise level above 100 kHz is approximately V/, limited by the digitizer sensitivity. Several narrow 1/ components appear between 140 and 320 kHz. At lower frequencies, the PSD rises due to voltage noise contributions shaped by the detector input capacitance (estimated at 4 pF, including the pogo-pin and front-end board input pads).
Beyond simply characterizing the ASIC gain behavior, this PSD analysis also enabled evaluation of different mechanical and electrical configurations of the system. We compared measurements with the back-end electronics housing open and closed, to evaluate how shielding the DC–DC converter present in the bias board is influencing the noise coupling. In addition, we also compared the configurations combining the use of indium foil pads and two different screws material (PEEK vs. stainless steel) used for sensor and front-end PCB interconnection. As shown in Fig. 7(b), the configuration with indium foil and INOX (Stainless steel) screws consistently resulted in lower noise levels across the three ASIC gain settings (HG, MG, LG).
@@ 3.3 Fluorescence Measurements with lab X-ray generator source
To characterize the detector response with varying photon flux, fluorescence spectra were acquired with samples of 55Mn, 63Cu, 119Sn, and 137Ba using a laboratory X-ray generator. The generator was equipped with a molybdenum (Mo) anode, operated at a maximum voltage of current of 50 kV and 40 mA, respectively. The maximum photon energy generated corresponds to the applied voltage, reaching up to 50 keV in this setup. A set of multiple spectra was recorded across a range of peaking times (PT) from 0.5 to 12 µs, with acquisition windows of 20–30 s per spectrum. The input count rate (ICR) was systematically varied from 6 kcps to 1.2 Mcps.

### v2

# Development and first performance evaluation of multi-element monolithic HPGe detector for X-ray spectroscopy

@@ Abstract
The first operational prototype of a high-purity Germanium (HPGe) detector developed within the European LEAPS-INNOV project is presented in this work. This prototype features a monolithic, multi-element sensor optimized for high-resolution X-ray spectroscopy in the hard X-ray regime, capable of handling high count rates (20-250 kcps/mm2) across a broad energy range (5-100 keV). We discuss here a complete laboratory-based characterization of the detector’s performance, as well as an on-beam evaluation at the BM05 beamline of the ESRF synchrotron facility, using monochromatic X-rays in the 20–50 keV energy range. We provide a detailed performance assessment that also includes a phenomenological defect-depth estimation model.
@@ 1 Physics Motivation
Fundamentally, the choice of Germanium over Silicon is motivated by its higher atomic number (Z=32) vs. (Z=14) for Si, which enhances the photoelectric absorption cross-section, particularly in the hard X-ray regime, meaning that HPGe detectors exhibit substantially higher quantum efficiency for photon energies above 20 keV [MOHANTY2008186], which allows an effective detection of weak signals in dilute samples containing heavy elements [Cole2001RareEarthEXAFS]. Additionally, Germanium’s smaller bandgap (0.67 eV) enables finer energy resolution by generating more charge carriers per absorbed photon, improving the resultant signal to background/noise ratio (SNR) [FARROW1995567, dennis2019].
@@
The monolithic design avoids boundaries between separate crystals, which can cause signal variations due to gaps or misalignment. Using high-purity germanium reduces crystal defects and, hence, reduces charge-carrier trapping. As a result, the detector achieves high energy resolution and good signal response across the entire active area [Hexgonal_pad_N_tartoni], thereby attaining qualities critical for spectroscopy applications.
@@
Great care is required during fabrication, as such detectors require ultra-high purity and operate in a cryogenic environment (near 77 K) to suppress thermal noise.
@@
Building on this motivation, Work Package 2 of the European LEAPS-INNOV [LEAPSINNOV2020] project has taken a step towards developing a new generation of HPGe detectors [F_ORSINI_FIRSTARTICLE, Goyal_2025, Gimenez_2025]. We have developed two such detectors with 10-element monolithic sensors, featuring pixel areas of 5 mm2 and 20 mm2. These detectors should offer the capability of handling high throughput in a broad X-ray energy range, while maintaining a reliable energy resolution aiming 180 eV FWHM at 5.9 keV, within the integration time window/Peaking Time (PT) of 1 s [Goyal:2025oph]. In addition, the Digital Pulse Processing technique (DPP) will further improve detector performance by electronically rejecting charge-sharing events, which are responsible for degrading the resultant spectral resolution [QuantumXspress3] due to incomplete charge collection.
@@
This article presents the full development and characterization of the first HPGe prototype. In Section 2, we describe the detector design, including the monolithic design of the sensor, the cryogenic cooling, mechanical architecture, and TETRA ASIC-based new front-end electronics. Section 3 discusses the laboratory-based characterization with detailed analysis of signal rise time, Power spectral density, fluorescence measurements with metallic foil samples, and surface scanning using a micro spot beam. In Section 4, we discuss the first on-beam evaluation of the detector, followed by a detailed analysis of defects observed in several pixels. We provide a phenomenological model to derive the depth of these defects. Finally, we present a summary of detector performance to date and outline future directions for performance improvement.
@@ 2 The XAFS-DET germanium detector
The XAFS-DET detector, shown in Figure 1, is composed of a compact detector head, a detector body, and ancillary elements. Other synchrotron spectroscopy detectors typically follow the same schematic structure. The detector head, which includes the germanium sensor, the front-end electronics, and their interconnections, is under vacuum at cold temperature (77 ). The detector body comprises the cooling system, the cryostat with its flanges, and the back-end electronics. Ancillary elements include the digital pulse processor, voltage power supplies, and equipment control and monitoring.
@@ 2.1 Sensor Design
The sensor is a multi-element high-purity germanium sensor manufactured by Mirion Technologies (CANBERRA). A scheme of the sensor is shown in Figure 2. It has a square shape (20 mm length) and a thickness of 4.1 mm. The sensor is a p-type germanium bulk operating in hole-collection mode, which is expected to reduce charge trapping relative to electron collection. The front side (facing the incident X-rays) is phosphor-implanted (n-type doped). Whereas, the back side is boron-implanted (p-type doped) and is segmented into a central hexagonal pixel, six trapezoidal pixels situated in a ring around the central one, and three outer ring-shaped pixels, which serve as ancillary pixels to reject charge-sharing events by sharing the incident charge with the guard ring. The dimensions of hexagonal and trapezoidal pixels are defined to achieve an area of 20 mm2 per pixel. Aluminium contacts were evaporated on both sides, and the inter-pixel spacing provides a simulated insulation value of 10 T. A positive potential difference is established by applying a high voltage to the frontside window contact via a high-voltage Rogers PCB, while backside pixels are connected to front-end board pads via copper-gold-plated pogo pin connectors (Model P706 from Peak Test, 6.9 mm long with 0.2 mm-diameter tips). The full depletion voltage is around +33 V, and we operated the detector in the range of +75 to +300 V during characterization. We measured the leakage current of each pixel to be less than 1 pA at +100V while operating at liquid nitrogen temperature.
@@ 2.2 Electronics
TETRA ASIC is a four-channel charge-sensitive (CSA) preamplifier working in hole collection mode with synchronous reset [6154396]. The preamplifier gain setting (low, medium, or high), which varies for feedback capacitances of 29, 58, and 174 fF, respectively, Reset features (threshold, width, and external trigger), and channel disabling can be configured by the user via a 32-bit shift register. All four channels share a test-pulse input to characterize signal fall time and crosstalk between channels without the sensor, as well as to simulate input leakage current [CMOS_CSA]. The FEB is a circular board (20 mm radius, 1.4 mm thick) made of ceramics, as shown in Figure 3 (left). It links the ten pogo pin pads (0.4 mm radius), on its front side, to ten input channels of three TETRA ASICs (two with three channels enabled, one with all four channels enabled) located on its back side. The FEB back side also houses a thermistor. Output signals and TETRA power supplies are routed to the BEB via a 60 cm-long flex kapton PCB cable with a Sub-D 25 connector. The cable is glued to the FEB back surface, and electrical continuity is achieved via wire bonding, while the Sub-D 25 connector is screwed into the back-flange feedthrough of the vacuum cube.
@@
The BEB comprises three PCB boards (buffer, bias, and control) designed to bias and configure three TETRA ASICs and buffer 10 analog channels. A conceptual schema is shown in Figure 3 (right). The buffer board is directly connected to the subd-25 feedthrough to minimize the path of ASIC outputs. The ten signals are buffered and routed to ten SMA connectors, which are connected to the Digital Pulse Processor via coaxial RF cables. The bias board supplies power to itself and the other two boards, applies voltage-level shifting to the ASIC program bus, and acquires the FEB thermistor signal. The bias board also includes three SMA connectors: two inputs for the external reset pulse and the test pulse, and one output for the inhibit reset signal. The control board monitors and controls BEB features via an FPGA unit and communicates with the user via USB 3.0 or Ethernet.
@@
The electronics performance is summarized in Table 1. The step function response without the sensor was characterized at room temperature (23-25°C) in terms of signal gain and fall-time (10% to 90% of the amplitude) for the three gain settings and for all channels. These features were measured from the mean waveform generated by a squared-test pulse (1 V amplitude, 10 ns rising edge), as illustrated in Figure 4(a). Fall-time values of 19-23 ns were measured for all gains, and we verified that the gain ratios high-medium and medium-low were 2.0 and 3.0, see Figure 4(b). We also quantified absolute gain values during subsequent laboratory tests using the sensor. In a second step, the range of values for reset duration (0.7-5.9 s) and threshold (-2.3-1.3 V) was measured. The reset period was longer than 4-5 sec, for a negligible leakage current. The period was shorter (around 23, 41, and 116 ms for high, medium, and low gain settings, respectively) with the leakage emulator enabled. Finally, the Equivalent noise charge (ENC) for peaking times between 0.1 and 10 s was measured, yielding values of 36-45 and 29-35 electrons for 1.0 and 10 s, respectively, within specifications. This series of tests was repeated at cold temperature (113 K), as discussed in [Goyal:2025oph].
@@
| ENC ( 1 s PT ) | < 45 electrons | 36-45 electrons |
| ENC ( 10 s PT ) | < 35 electrons | 29-35 electrons |
@@ 2.3 Cooling and mechanical design
The aluminum container, a copper arm holder, a horizontal copper cooling pipe, a thermal strap, and a vertical copper braid to the cryocooler evacuate the dissipated power from the front-end board [Quispe:2024wis].The pads of Indium foils link each interface. The cooling pipe is supported by two PEEK arm holders inside the inner cylinder, with small supports to minimize thermal losses. Two thermometers, glued to the vertical copper braid and the copper arm, monitor the temperature. The second one works as a feedback device for the cryocooler system. A zeolite box is attached to the vertical copper braid piece, which maintains vacuum integrity in static vacuum by adsorbing residual gases using pre-activated zeolite treated under vacuum at high temperatures. The box is equipped with a heating element and a thermometer, which are used to regenerate zeolites once they are saturated. The cooling chain of the detector is showcased using Figure5 .
The cryocooler is mounted on a vacuum cube that holds the detector head in a front-facing tube. The cube’s three CF100 flanges provide dedicated feedthroughs for electronics (Sub-D 25), cooling thermometers (SubD9), and HV bias. A separate flange houses the pumping port, vacuum gauge, and a SubD9 for the zeolite heater. A low-pass RC filter with a 6.5 Hz cut-off frequency conditions the HV signal to the sensor. A turbo pumping system achieves a vacuum of mbar, with residual gas analysis (MKS model) confirming only trace amounts of air and water.
Induced vibrations at the sensor head were measured using an Active Vibration Cancellation (AVC) system (model VSM101) produced by ifm Electronics. Acceleration values of 0.04-0.07 m/s2 at 1-2 frequencies above 50 Hz were measured in the three axes, which are within specifications.
@@ 3 Detector characterization in the laboratory
The procedures and results for each test are discussed in detailed in the following subsection.
@@ 3.1 Mean Signal risetime and preamplifier gain
The signal rise time and preamplifier gain were characterized using a 55Fe radioactive source emitting 5.9 keV and 6.4 keV X-rays. We operate the detector under low-flux conditions (approximately 10 kcps), and subsequently we acquire 105 individual signal waveforms using a Lecroy digital oscilloscope operating at 0.5 Gs/s. We performed the measurements with bias voltages varying from +50 V to +330 V in 10 V increments.
Each recorded waveform was processed to extract the signal amplitude and rise time. The distribution of these parameters is analyzed as a function of bias voltage. The mean pulse amplitude was used to determine the preamplifier gain, yielding values between 1.99 and 2.06 mV/keV with a systematic uncertainty of 0.02 mV/keV, which are within design specifications.
As shown in Figure 6, the rise time distributions shift progressively toward shorter values as the bias voltage increases. This trend reflects the expected enhancement in hole drift velocity with stronger electric fields, reaching a minimum rise time of about 42 ns in the 280–330 V range.
@@ 3.2 Power Spectral Density Measurements
We measured the PSD over the frequency range of 100 Hz to 50 MHz using a fast digitizer (Model DT9862, by Data Translation, triggered 100 ms after the RESET signal to capture noise during the signal-processing window. Measurements were performed on the central pixel across all available TETRA ASIC gain settings. As shown in Figure 7(a), the noise level above 100 kHz is approximately V/, limited by the digitizer sensitivity. Several narrow 1/ components appear between 140 and 320 kHz. At lower frequencies, the PSD rises due to voltage noise contributions shaped by the detector input capacitance (estimated at 4 pF, including the pogo-pin and front-end board input pads).
Beyond simply characterizing the ASIC gain behavior, this analysis also enabled evaluation of different mechanical and electrical configurations of the system. We compared measurements with the back-end electronics housing open and closed, to evaluate how shielding the DC–DC converter present in the bias board is influencing the noise coupling. In addition, we also compared the configurations combining the use of indium foil pads and two different screws material (PEEK vs. stainless steel) used for sensor and front-end PCB interconnection. As shown in Figure 7(b), the configuration with indium foil and INOX (Stainless steel) screws consistently resulted in lower noise levels across three gain configurations.
@@ 3.3 Fluorescence Measurements with lab X-ray generator source
To characterize the detector response as a function of photon flux, fluorescence spectra were acquired with samples of 55Mn, 63Cu, 119Sn, and 137Ba using a laboratory X-ray generator. The generator was equipped with a molybdenum (Mo) anode, operated at a maximum voltage current of 50 kV and 40 mA, respectively. The maximum photon energy generated is determined by the applied voltage, reaching up to 50 keV in this setup. A set of multiple spectra was recorded across a range of peaking times (PT) from 0.5 to 12 s, with acquisition windows of 20–30 s per spectrum. The input count rate (ICR) was systematically varied from 6 kcps to 1.2 Mcps.